# Coloring Grids Avoiding Bicolored Paths

### v4

@@ 2. Main Result
If j=k3j=k-3, then $C_{j+2}$ does not exist. In this case, it is not guaranteed to have a bicolored $P_{k}$ by the neighbors of $V(P)$ and itself. So, we use Lemma 2 to show the existence of a bicolored $P_{k}$. Let $C$ be the bicolored red-blue component containing $P.$ If $C$ contains any vertex from the top or bottom sides outside $P,$ then $C$ contains a red-blue $P_{k}.$ If $C$ contains any vertex from the left side outside $P,$ then the red-blue path in $C$ connecting $P$ to the left side together with $P$ yield a bicolored path with at least $k$ vertices. Thus, we assume that $C$ is not connected to any side of the grid (other than the right side) outside $P$. This implies that there exists a partial walk of $C$ from $x_{k-2}$ to $w_{1}$. However, Lemma 2 yields that $x_{k-2}$ and $w_{1}$ have the same color, a contradiction.
Case 2.2. $i\equiv k$ (mod 2) and $i$, even.
[…]
Case 2.3. $i\not\equiv k$ (mod 2) and i, odd.
If $j\not=1$, by Remark 4, the union of the two red-green paths $v_{1},v_{2},\dots,v_{i}$ and $w_{i},\dots,w_{k-2}$, together with $x_{k-2}$ induce a bicolored $P_{k}.$ If $j=1$, then $C_{j+2}$ exists. By Remark 4, $x_{1},\dots,x_{i-1}$ and $y_{i},\dots,y_{k-2}$ are red-green paths. If $y_{i-1}$ (black vertex in Figure 4 (c)) has color blue, then there is a red-blue $P_{k}$ induced by $\{w_{1},\dots,w_{i-1},x_{i-1},y_{i-1},y_{i},x_{i},$ $\dots,x_{k-2}\}.$ Otherwise, $y_{i-1}$ has color green and there is a red-green $P_{k}$ induced by $\{w_{1},x_{1},\dots x_{i-1},$ $y_{i-1},\dots,y_{k-2},x_{k-2}\}.$
Case 2.4. $i\equiv k$ (mod 2) and $i$, odd.
There is a red-blue $P_{k}$ induced by $\{w_{1},\dots,w_{i-1},x_{i-1},x_{i},w_{i},w_{i+1},x_{i+1},\dots,x_{k+2}\}.$
@@
As a special case, if $i=1$, then $P$ crosses between columns at the top row having vertex set $\{w_{1},x_{1},x_{2},\dots,x_{k-2}\}$. By Remark 4, $w_{2},\dots,w_{k-2}$ induce a red-green path and $k$ must be even, since $w_{k-2}$ is green by maximality of $P$. If $C_{j-1}$ exists, $v_{1}$ can be either green or blue in a proper coloring, inducing either a red-green $P_{k}$ together with $w_{1},\dots,w_{k-2},x_{k-2}$ or a red-blue $P_{k}$ together with $P$, respectively. In case, $j=1$, then $C_{j+2}$ exists and the path $y_{2},\dots,y_{k-2}$ is a red-green colored path by Remark 4. The red-green colored paths on $C_{j}$ and $C_{j+2}$ together $x_{k-2}$ yield a red-green $P_{k}.$ The same proof holds also for i=k2i=k-2 by symmetry.
@@
Let $P_{k-2}\square P_{k-2}$ be colored with colors $\{c_{1},c_{2},c_{3}\}$ and assume that such a 3-coloring does not have a peripheral bicolored component. Let $C$ be a partial, $c_{1}$-$c_{2}$ colored component containing vertices from the top side and possibly left side of the grid. Let $B_{C}=(v_{1},v_{2},\dots,v_{r})$ be a partial walk of $C$. We need the following definitions associated with $C$:
  • $D_{C}$: By Lemma 2, there is a $c_{1}$-$c_{3}$ colored connected subgraph, $c_{1}$ being the color of $v_{1}$, induced by $v_{i}$ and $u_{(i+1)/2}$, for odd $i$, $1\leq i\leq r-2,$ where $u_{(i+1)/2}\not\in C$ is the green vertex on the $C_{4}$ containing $\{v_{i},v_{i+1},v_{i+2}\}.$ Let $D_{C}$ be the bicolored component containing this $c_{1}$-$c_{3}$ colored connected subgraph.
  • $\mathcal{B}_{C}$: Let this be the collection of $C_{4}$’s induced by $\{u_{(i+1)/2},v_{i},v_{i+1},v_{i+2}\}$, for odd $i$, $1\leq i\leq r-2$. As $r\geq 3$, $\mathcal{B}_{C}\neq\emptyset$.
@@
    Both $C$ and $D_{C}$ have some vertices from the top and left side of the grid. Thus, $v_{1}$, $v_{1}^{\prime}$ are on the top side of the grid, $v_{r}$, $v_{r}^{\prime}$ are on the left side of the grid.
@@
Note that, after an iteration, the squares enclosed by the $C_{4}$’s in $\mathcal{B}_{C}$ are missing in $R_{C}$. As $\mathcal{B}_{C}\neq\emptyset$, $R_{C}$ becomes smaller after each iteration in cases (1)-(3). Thus, as long as $C$ is not peripheral, $v_{1}=v_{1}^{\prime}$ or $v_{r}=v_{r}^{\prime}$ can happen at most once for each vertex, as there are exactly two $C_{4}$’s that contain $v_{1}$ (same for $v_{r}$) and no $C_{4}$ belongs to $R_{C}$ in more than one iteration. In that sense, each vertex plays the role of $v_{1}$ or $v_{r}$ with multiplicity at most two throughout all iterations. Consequently, at each consecutive pair of iterations, the position of $v_{1}$ approaches the top right corner by at least one edge and the position of $v_{r}$, approaches the top left or the bottom left corner of the grid by at least one edge, depending on the case. These observations show that after some finite number of iterations, none of the cases (1)-(3) will hold and $D_{C}$ becomes a peripheral bicolored component. ∎

### v5

@@ 2. Main Result
If j=k3j=k-3, then $C_{j+2}$ does not exist. In this case, it is not guaranteed to have a bicolored $P_{k}$ only using vertices in $P$ and their neighbors. So, we use Lemma 2 to show the existence of a bicolored $P_{k}$. Let $C$ be the bicolored red-blue component containing $P.$ If $C$ contains any vertex from the top or bottom sides outside $P,$ then $C$ contains a red-blue $P_{k}.$ If $C$ contains any vertex from the left side, then the red-blue path in $C$ connecting $P$ to the left side together with $P$ yield a bicolored path with at least $k$ vertices. Thus, we assume that $C$ is not connected to any side of the grid (other than the right side) outside $P$. This implies that there exists a partial walk of $C$ from $x_{k-2}$ to $w_{1}$. However, (4) of Lemma 2 yields that $x_{k-2}$ and $w_{1}$ have the same color, a contradiction.
@@
As a special case, if $i=1$, then $P$ is $w_{1},x_{1},x_{2},\dots,x_{k-2}$ and crosses between columns at the top row. By Claim 4, $w_{2},\dots,w_{k-2}$ is a red-green path and $k$ must be even, since $w_{k-2}$ is green by maximality of $P$. If $C_{j-1}$ exists, $v_{1}$ can be either green or blue in a proper coloring, inducing either a red-green $P_{k}$ together with $w_{1},\dots,w_{k-2},x_{k-2}$ or a red-blue $P_{k}$ together with $P$, respectively. In case, $j=1$, $C_{j+2}$ exists and the path $y_{2},\dots,y_{k-2}$ is a red-green colored path by Claim 4. The red-green colored paths on $C_{j}$ and $C_{j+2}$ together $x_{k-2}$ yield a red-green $P_{k}.$
@@
Let $P_{k-2}\square P_{k-2}$ be colored with colors $\{c_{1},c_{2},c_{3}\}$ and assume that such a 3-coloring does not have a peripheral bicolored component. Let $C$ be a partial, $c_{1}$-$c_{2}$ colored component containing vertices from the top side and possibly left side of the grid. Let $B_{C}=(v_{1},v_{2},\dots,v_{r})$ be a partial walk of $C$, where $v_{1}$ has color $c_{1}$. We need the following definitions associated with $C$:
  • $D_{C}$: By Lemma 2, there is a $c_{1}$-$c_{3}$ colored connected subgraph, induced by $v_{i}$ and $u_{(i+1)/2}$, for odd $i$, $1\leq i\leq r-2,$ where $u_{(i+1)/2}\not\in C$ is the $c_{3}$-colored vertex on the $C_{4}$ containing $\{v_{i},v_{i+1},v_{i+2}\}.$ Let $D_{C}$ be the bicolored component containing this $c_{1}$-$c_{3}$ colored connected subgraph.
  • $\mathcal{W}_{C}$: Let this be the collection of $C_{4}$’s induced by $\{u_{(i+1)/2},v_{i},v_{i+1},v_{i+2}\}$, for odd $i$, $1\leq i\leq r-2$. As $r\geq 3$, $\mathcal{W}_{C}\neq\emptyset$.
@@
    Both $C$ and $D_{C}$ have some vertices from the top and left side of the grid. Thus, $v_{1}$, $v_{1}^{\prime}$ are on the top side of the grid, $v_{r}$, and $v_{r}^{\prime}$ are on the left side of the grid.
[…]
[width=0.5]main_iteration.jpg
Figure 5. An example according to the scenario (2), the left and right grids showing the current $C$ before and after the iteration (and the flip), respectively. $\mathcal{R}_{C}$ is shown with the shaded region and $B_{C}$ is marked with bold edges.
@@
Note that, after an iteration, the squares enclosed by the $C_{4}$’s in $\mathcal{W}_{C}$ are missing in $R_{C}$. As $\mathcal{W}_{C}\neq\emptyset$, $R_{C}$ becomes smaller after each iteration in cases (1)-(3). Thus, as long as $C$ is not peripheral, $v_{1}=v_{1}^{\prime}$ or $v_{r}=v_{r}^{\prime}$ can happen at most once for each vertex, as there are exactly two $C_{4}$’s that contain $v_{1}$ (same for $v_{r}$) and no $C_{4}$ belongs to $R_{C}$ in more than one iteration. In that sense, each vertex plays the role of $v_{1}$ or $v_{r}$ with multiplicity at most two throughout all iterations. Consequently, at each consecutive pair of iterations, the position of $v_{1}$ approaches the top right corner by at least one edge and the position of $v_{r}$, approaches the top left or the bottom left corner of the grid by at least one edge, depending on the case. These observations show that after some finite number of iterations, none of the cases (1)-(3) will hold and $D_{C}$ becomes a peripheral bicolored component. ∎